Y=θ [SIN(INθ)+COS(INθ)],THEN FIND dy÷dθ.

Y=θ[SIN(INθ)+COS(INθ)],THEN FIND dy÷dθ.

Grade:10

2 Answers

Yogita Bang
11 years ago

y=Θ[sin(lnΘ)+cos(lnΘ)]

dy/dΘ=[sin(lnΘ)+cos(lnΘ)] + Θ[cos(lnΘ)-sin(lnΘ)]*1/Θ ---->(Use Multiplication rule)

dy/dΘ=2cosΘ

 

If u like answer please click YES below Smile

Rohan Das
11 years ago

 WE KNOW

SIN(INθ)+COS(INθ)=PIE/2

 

y=θ*pie/2

dy÷dθ=pie/2

so